OpenStudy (anonymous):

Factor. m2 + 12m + 36 (m + 6)^2 (m – 6)^2 Cannot be factored

OpenStudy (anonymous):

and (m + 36)^2

OpenStudy (irishboy123):

Clues 6*6 = 36 6+6 = 12 But 6*-6 = -36 6-6= 0 Clear?